A sinusoidal wave of frequency f=100 hz propagates in a medium with a speed 12m/s.Give the expressions of the distance separating:
A) two in-phase points
B) two anti-phase points
1
Expert's answer
2019-10-02T09:38:25-0400
"\lambda=\frac{v}{f}=\frac{12}{100}=0.12\ m"

a)


"x=n\lambda=(0.12n)\ m"

b)


"x=(n+0.5)\lambda=(0.12n+0.06)\ m"